OpenStudy (savy):

0.35=1-293.15/x please help me solve for x

OpenStudy (savy):

\[0.35=1-\frac{ 293.15 }{ x }\]

OpenStudy (deeloucook):

x=451 That's the answer i just got

OpenStudy (savy):

how please?

563blackghost (563blackghost):

First we must rid of the decimal and 1 in order to do that we multiply by 100. \(\LARGE\bf{0.35 \times 100 = 1 \times 100 - \frac{293.15}{x} \times 100}\) This resorts to. \(\huge\bf\color{purple}{35=100-\frac{29315}{x}}\) We then multiply both sides by x. \(\huge\bf\color{orange}{35x=100x-29315}\) In which we would subtract `100x` from both sides... \(\huge\bf\color{green}{-65x=-29315}\) Now we would divide by `-65`. \(\huge\bf{\frac{-65x}{-65}=\frac{-29315}{-65}} \rightarrow \color{red}{x = 451}\)

563blackghost (563blackghost):

That should do it ;)

563blackghost (563blackghost):

@deeloucook you can not be giving the answers away, you must help the owlet so that they may understand the question and how to solve. This is in violation of the code of conduct. Please refrain from this in the future.

OpenStudy (savy):

thank you @563blackghost